### v1

# QOS: A Quantum Operating System

Abstract.
We introduce the Quantum Operating System (QOS), a unified system stack for managing quantum resources while mitigating their inherent limitations, namely their limited and noisy qubits, (temporal and spatial) heterogeneities, and load imbalance. QOS features the QOS compiler—a modular and composable compiler for analyzing and optimizing quantum applications to run on small and noisy quantum devices with high performance and configurable overheads. For scalable execution of the optimized applications, we propose the QOS runtime—an efficient quantum resource management system that multi-programs and schedules the applications across space and time while achieving high system utilization, low waiting times, and high-quality results.
We evaluate QOS on real quantum devices hosted by IBM, using 7000 real quantum runs of more than 70.000 benchmark instances. We show that the QOS compiler achieves 2.6–456.5$\times$ higher quality results, while the QOS runtime further improves the quality by 1.15–9.6$\times$ and reduces the waiting times by up to 5$\times$ while sacrificing only 1–3% of results quality (or fidelity).
1. Introduction
Quantum Cloud Computing. Quantum computing promises to solve computationally intractable problems with classical computers (Arute et al., 2019; Farhi et al., 2014). Thanks to remarkable technological advances in materials science and engineering (Siddiqi, 2021; Hanneke et al., 2010), quantum hardware has become a reality in the form of quantum processing units (QPUs) that consist of quantum bits (qubits) (ibmQuantum, [n. d.]). Interestingly, QPUs are now readily available in a quantum-as-a-service fashion offered by all major cloud providers (ibmQuantum, [n. d.]; awsQuantum, [n. d.]; googleQuantum, [n. d.]; azurequantum, [n. d.]).
While quantum hardware is now a reality, the associated quantum software systems are rudimentary. These QPUs face classical OS challenges that our systems community has tackled in the past, including scalability, performance, efficiency, faults (a.k.a. errors), scheduling, and utilization (Bravyi et al., 2022). Unfortunately, no operating system exists to tackle these challenges holistically for modern quantum hardware.
Fundamental Challenges of QPUs. A natural tendency would be to treat these QPUs as yet another accelerator class (e.g., GPU, TPU, FPGA) and manage them as accelerator as-a-service to offload compute-intensive tasks. We argue that this approach might be sub-optimal or even flawed!
The reason is that QPUs present fundamentally unique hardware-level challenges that the systems community has not considered and cannot be directly mapped to classical accelerator-oriented computing (we empirically detail these hardware challenges in §$\S$ 3). In particular, QPUs operate in the NISQ-fashion (Noisy Intermediate-Scale Quantum (Preskill, 2018)), leading to a non-deterministic computing platform, where even two QPUs with identical qubits exhibit completely different behaviors across space and time (Ravi et al., 2021a; Patel et al., 2020).
More specifically, QPUs are inherently noisy and small in computational capacity (Preskill, 2018), which limits the size of the problems they can solve. Second, the degree of noise differs across QPUs, even of identical architecture and model, making it difficult to decide which QPUs should execute a quantum program without compromising performance (Ravi et al., 2022b). In addition, we can not trivially multi-program multiple quantum programs on the same QPU to increase utilization since QPU qubits can interfere with each other in undesirable and unpredictable ways (Murali et al., 2020), severely degrading performance (Liu and Dou, 2021). Finally, it is generally impossible to save or copy a quantum program during execution (Park, 1970), which further limits scheduling opportunities for preemption or resource sharing in general.
State-of-the-Art of Quantum Software Systems. The current state of software can be roughly compared to IBM mainframe batch OSes from the 60s, where the QPUs are managed through rudimentary interfaces. Researchers have proposed specialized approaches to address some of the aforementioned OS and QPU challenges individually, for instance, performance (Tang et al., 2021a), multi-programming (Das et al., 2019), or scheduling (Ravi et al., 2021b). Unfortunately, these proposed approaches are designed to solve an individual issue, which prevents them from being composed together or with other OS mechanisms to create a holistic software stack. To leverage quantum computing practically, we must address the key challenge of combining such mechanisms in a unified software stack for quantum computing.
Novelty. However, designing a unified system stack that supports general OS abstractions while addressing the QPU challenges is not trivial. The system should support cross-stack software mechanisms, from the compiler level for quantum program optimization to the runtime level for QPU resource management. More specifically, we require a modular and extensible compiler infrastructure for increasing execution quality, multi-programming for increased utilization, and scheduling for load balancing, all in the presence of QPU noise and heterogeneities. This way, the system can achieve the cloud users’ goals, i.e., high-quality quantum computation and low waiting times, and the quantum cloud operator’s goals, i.e., QPU resource efficiency and scalability.
Figure 1. Example of a typical quantum algorithm (§ 2.1) (a) Input graph for max-cut. (b) The quantum circuit encoding the formulation of max-cut for the graph. (c) The result of circuit execution is a probability distribution of bitstrings. (d) The result of (c) is interpreted as a max-cut between vertices $\{0,1,4\}$ and $\{2,3,5\}$.
QOS: A Unified System Stack for Quantum Computing. We propose QOS, an end-to-end system for holistically tackling quantum computing challenges. QOS provides a unified architecture for supporting compiler and OS mechanisms with pluggable and configurable policies. In QOS, we implement such policies to achieve the aforementioned users’ and operator’s goals. To achieve this, QOS builds on a unified abstraction and comprises two main components:
  • The Qernel Abstraction: We introduce the Qernel abstraction that acts as a common denominator for the QOS mechanisms to apply their policies (§ 5.1). A Qernel contains the Qernel intermediate representation (QIR) and static and dynamic properties, leveraged by the QOS components to apply their policies.
  • QOS Compiler: We introduce the QOS compiler (§ 5, 6, 7), an extensible and modular compiler workflow that leverages the QIR and static properties to optimize quantum programs for increased execution quality.
  • QOS Runtime: We present the QOS runtime (§ 8), a scalable system for QPU resource efficiency. The system offers automated QPU selection to abstract heterogeneity away, multi-programming to increase QPU utilization, and load-aware scheduling to achieve low waiting times while maintaining high execution quality.
Figure 2. Technical Foundations (§ 2.2) (a) The quantum circuit of Figure 1. (b) The physical layout of an IBM Falcon QPU. (c) The transpiled circuit with the QPU’s noise sources.
We implement QOS in Python by building on the Qiskit framework (Qiskit contributors, 2023). We evaluate QOS on IBM’s 27-qubit QPUs (ibmQuantum, [n. d.]), using a dataset of more than 7000 quantum runs and 70.000 state-of-the-art quantum benchmark instances used in popular quantum algorithms (Tomesh et al., 2022; Quetschlich et al., 2022; Li et al., 2021). Our evaluation shows that the QOS compiler improves the quantum program properties by 51% on average, which leads to 2.6–456.5$\times$ improvement in the quality of the results, depending on the problem size (§ 9.2). The QOS runtime increases the quality of the results by 1.15–9.6$\times$ for the same target utilization (§ 9.4) and reduces the waiting times by 5$\times$ while sacrificing at most 3% of the quality of the results (§ 9.5).
2. Background
2.1. Quantum Computing 101: An Example
Let us understand the basics of quantum computing using the classic max-cut problem. This simple combinatorial optimization problem is expressed in the quantum world as the Quantum Approximate Optimization Algorithm (QAOA) (Farhi et al., 2014). Figure 1 shows a high-level example of how QAOA solves a max-cut problem of the input graph of (a). To solve it, the problem must be first encoded as a quantum circuit (Figure 1 (b)), which consists of quantum bits (qubits) and quantum gates that exhibit quantum mechanical properties. Here, we use as many qubits as the number of nodes of the input graph, where each qubit $q_{i}$ corresponds to a graph vertex $i$. To change the state of the qubits, we apply quantum gates over time, from left to right. There are two types of gates: 1-qubit gates (e.g., NOT gate) and 2-qubit gates (e.g., XOR gate). Finally, at the end of the circuit, we measure each qubit to read its value (0 or 1), which gives bitstrings as output.
Unlike classical circuits, which operate deterministically, quantum circuits are inherently probabilistic. The reason is that qubits exhibit quantum mechanical properties, such as superposition. In the superposition state, the qubit is not 0 or 1, but it is both simultaneously (recall Schrodinger’s cat experiment (schrondinger, [n. d.])). Therefore, quantum gates also have probabilistic effects; we can’t know the result until the final measurements (i.e., open the box and check the cat’s state). To obtain a meaningful result, we execute the circuit in many trials (“shots”), with each trial providing a specific bitstring from the qubit measurement. The solution of the quantum calculation is, therefore, a probability distribution over all possible bitstrings of the measured qubits (Figure 1 (c)).
Figure 3. (a) Challenge #1, Fidelity (§ 3.1). Impact of the number of qubits (circuit size) on fidelity. There is an average 98.9% reduction in fidelity from 4 to 24 qubits. (b) Challenge #2, Spatial heterogeneity (§ 3.2) Fidelity of a 12-qubit GHZ circuit on different IBM QPUs. There is a 38% fidelity difference from best to worst QPU.
In our example, the result of the final execution of the quantum circuit gives a probability distribution that represents the solutions of the max-cut problem. High probability maps to the solution, while low ($\sim 0$) does not represent a solution. Figure 1 (d) shows a solution for our example. It corresponds to the bitstring with the highest probability, 110010, which means that we have measured $1$ for the qubits $q_{0}$, $q_{1}$, and $q_{4}$; therefore, a partition contains vertices $\{0,1,4\}$.
2.2. Technical Foundations
Execution Model. The technology and engineering required to build QPUs renders them an expensive resource, thus, QPUs are mainly offered in the cloud as a quantum-as-a-service model (ibmQuantum, [n. d.]; googleQuantum, [n. d.]; awsQuantum, [n. d.]). To run quantum programs, users typically write circuit-level code (Figure 2 (a)), which then transpile on the QPU to make it executable, send it to the cloud for execution, and finally get the results back. Specifically, the transpilation process performs three key steps: (1) converting the gates of the circuit to the native gate set of the QPU, (2) mapping the logical qubits of the circuit to the physical qubits of the QPU, (3) routing the qubits to the physical qubits with restrictive connectivity by inserting SWAP gates. Figure 2 (b) shows the physical layout of an IBM Falcon QPU. Vertices are the physical qubits, and the edges capture their connectivity, i.e., between which qubits we can apply 2-qubit gates. Figure 2 (c) shows the physical circuit after transpilation with the QPU’s noise characteristics, which we detail next.
QPU Characteristics. Today’s QPUs are described as noisy intermediate-scale quantum (NISQ) devices (Preskill, 2018) since they exhibit low qubit numbers (e.g., up to a few 100s (ibmQuantum, [n. d.])) and are susceptible to hardware and environmental noise. Specifically, when measuring a qubit, there is a chance to read the opposite value, and when applying gates, there is a chance the gate performs a wrong operation (google-nisq-properties, [n. d.]). On top of that, when qubits are left idle (no gates applied) for more than a few hundred microseconds, the superposition decoheres to the |$\ket{0}$ state (Klimov et al., 2018), similar to resetting a register to 00. Lastly, qubits destructively interfere with each other via crosstalk effects (Cross et al., 2019). Figure 2 (c) shows qubits $Q_{2}$ and $Q_{3}$ that influence each other via crosstalk, noisy gates, qubit $Q_{5}$ that is left idle for long enough to decohere, and noisy measurements.
QPU Heterogeneity. Additionally, QPUs are vastly heterogeneous across space and time, unlike classical accelerators. Across space, QPUs vary in terms of technology, e.g., superconducting qubits (googleQuantum, [n. d.]; ibmQuantum, [n. d.]) or trapped ions (Häffner et al., 2008), architectures of the same technology, e.g., Falcon or Osprey superconducting QPUs (ibmQuantum, [n. d.]), and noise properties even for the same architecture (google-nisq-properties, [n. d.]), e.g., two identical QPUs exhibit different noise errors, etc. Across time, the QPUs are calibrated regularly to maintain their performance (Wittler et al., 2021; Tornow et al., 2022; ibmq-calibration, [n. d.]), a process that generates calibration data. These data quantify the noise errors, and change after each calibration cycle unpredictably.
Execution Quality. Lastly, to measure the quality of a circuit execution on NISQ QPUs, we use the fidelity metric (fidelity-qiskit, [n. d.]), which measures the similarity between the noisy probability distribution and the ideal probability distribution that noiseless, ideal QPUs can obtain. Fidelity is a number in the $[0,1]$ range, where a higher fidelity means a better quality result.
3. Motivation and Key Ideas
To motivate QOS, we present a set of unique challenges that distinguish QPUs from classical accelerators. We categorize our findings into four challenges that must be addressed to improve the practicality of quantum computing: fidelity, utilization, spatial and temporal heterogeneities, and load imbalance. The experimental methodology used is the same for the final system evaluation and is explained in detail in § 9.1.
3.1. Fidelity
Executing quantum programs with high fidelity is challenging since QPUs are characterized by relatively small numbers of qubits and noise, which leads to computation errors (§ 2.2). As the number of qubits and gates in a quantum circuit increases, the noise errors accumulate and the overall fidelity decreases.
Figure 4. (a) Challenge #2, Temporal variance (§ 3.2) Fidelity of a 6-qubit GHZ circuit on IBM Perth, across 120 calibration days. There are 20 pairs of days with more than 5% difference in fidelity. (b) Challenge #3, Utilization (§ 3.3) Maximum utilization achieved on a 27-qubit QPU for nine benchmarks while maintaining at least 0.75 fidelity. The average utilization is $26.3\%$, and the max is $29.6\%$. (c) Challenge #4, QPU Load (§ 3.4) Number of pending jobs on different IBM QPUs. The groups separated by vertical red lines indicate QPUs of the same size. There is up to 57$\times$ difference in number of jobs between QPUs of the same size.
Results. Our results are highlighted in Figure 3 (a). The x axis depicts the circuit size as the number of qubits while the y axis shows the fidelity, where higher is better. The experiment is run on the IBM Kolkata 27-qubit QPU. For each increase in qubits, the average fidelity decreases, up to $98.9\%$ from 4 to 24 qubits. Moreover, it is physically impossible to run circuits with a size larger than 27 qubits, since we cannot map them.
Implication. NISQ devices are limited due to size and noise and, therefore, cannot be practically used for large quantum circuits; either logically, because the circuit doesn’t fit in the device, or the execution results would be convoluted from noise errors, which translates to low fidelity.
Key Idea #1: Circuit Optimizations: To increase fidelity, we need a generic optimization infrastructure that transforms circuits into a physically and practically executable size.
3.2. Spatial and Temporal Heterogeneity
In the classical domain, two identical CPUs perform similarly for all applications, and at each point in time. In contrast, QPUs exhibit differences in the layout and connectivity of qubits (Gyongyosi and Imre, 2019) and variations in noise errors even for QPUs of the same model, which leads to spatial performance variance. Moreover, QPUs are calibrated regularly (§ 2.2), and after each calibration, the noise properties change (Wittler et al., 2021). As a result, the execution fidelity can vary across different calibration cycles, leading to temporal performance variance.
Results. Figure 3 (b) shows a 12-qubit GHZ circuit’s fidelity on different IBM QPUs. Fidelity varies across the QPUs, with a maximum difference of 38% from best to worst. Note that all six QPUs are of the same model (Falcon r5.11).
Figure 4 (a) shows a 6-qubit GHZ circuit’s fidelity over 120 calibration days executed on the IBM Perth 7-qubit QPU, where each data point represents a single day’s fidelity. The largest single-day difference in fidelity is 96.5%, and there are 20 instances of a single-day fidelity drop of more than 5%. Note that there is no way of predicting a QPU’s future calibration data to expect such performance differences.
Implications. Due to structural differences across QPUs, quantum circuits perform differently across them. Additionally, there is a high degree of temporal performance variance across calibration cycles, as the fidelity might change significantly from day to day with no discernible pattern.
Key Idea #2: Performance Estimation: We estimate a circuit’s potential performance on the available QPUs to automatically select the best-performing candidate(s).
3.3. Utilization
The fidelity of circuits decreases as their size increases (§ 3.1), and as a result, it becomes more challenging to utilize a QPU effectively. In contrast to the classical domain, where a CPU can be fully utilized, to get high-fidelity results in the quantum domain, we necessarily under-utilize QPUs.
Results. Figure 4 (b) shows the maximum utilization of the IBM Kolkata 27-qubit QPU for nine benchmarks while maintaining at least 0.75 fidelity. No benchmark exceeds 30% utilization, while the average is 26.3%. Higher fidelity values would yield even lower utilization and vice-versa.
Implications. There is a tradeoff between QPU utilization and performance (fidelity). In general, the lower utilization, the higher fidelity, and vice-versa. In contrast to the classical domain, the tension between these objectives is vastly larger.
Key Idea #3: Multi-programming: We spatially multiplex quantum circuits to increase system utilization (also known as multi-programming (Das et al., 2019)), and when combined with circuit optimizations, it also increases fidelity.
3.4. QPU Load Imbalance
The quantum cloud faces QPU load imbalance. The root cause is spatiotemporal heterogeneity (§ 3.2), combined with the manual QPU selection offered by the current quantum cloud model (ibmQuantum, [n. d.]). This leads to users selecting the “best performant” QPU based on empirical or arbitrary metrics (Ravi et al., 2021a).
Results. Figure 4 (c) shows the average number of pending jobs for different IBM QPUs across October 2023. The groups of QPUs (separated by the red dashed line) have a size of 7, 27, and 127 qubits, respectively. There is a 49$\times$, 57$\times$, and 5.7$\times$ maximum load difference across the groups, respectively.
Implications. Load imbalance leads to long waiting times for the users and thus, low quality of service. Additionally, there is no 1-1 mapping between the load and performance differences between QPUs. For instance, the 12-qubit GHZ circuit in Figure 3 (b) performs 1.1$\times$ better on IBM Hanoi than IBM Cairo, yet the former exhibits 57$\times$ higher load.
Key Idea #4: Load-aware Scheduling: We schedule (temporally multiplex) quantum circuits in a load-aware manner to balance the tradeoff between fidelity and waiting times.
4. Overview
Quantum computing is characterized by four main challenges that limit its practicality: (1) Execution fidelity is hindered by the small and noisy QPUs. (2) In contrast to classical accelerators, QPUs exhibit vast spatiotemporal heterogeneities, which renders their performance non-deterministic in both dimensions. (3) QPUs are heavily underutilized to give high-fidelity results. (4) QPUs face vast load imbalance, which leads to prolonged waiting times for the users.
Existing work is narrow-scoped and focuses on tackling one challenge at a time, but unfortunately, there are two main issues with this point solution approach. Firstly, composing the individual mechanisms to address all challenges at once is impossible without a common and unified infrastructure. Secondly, without synergies between the individual mechanisms, it is hard to maximize the objectives of the users, i.e., high fidelity and low waiting times, and the objectives of the quantum cloud operator, i.e., resource efficiency.
To this end, we propose QOS, an end-to-end system that tackles the challenges of quantum computing holistically. QOS strives for three design goals: (1) A unified architecture that supports compiler and OS mechanisms with pluggable policies and tunable configuration for managing the tradeoffs of QC. (2) QOS should enable the execution of large quantum circuits with high fidelity and scale with increasing incoming workloads and additional QPUs. (3) QOS should be resource efficient by achieving high QPU utilization and balancing QPU load to minimize waiting times.
Figure 5. QOS overview (§ 4): QOS consists of two main components: the QOS compiler (top) and the QOS runtime (bottom). Below QOS lie the QPU devices and classical nodes.
4.1. The QOS Architecture
Figure 5 shows the overview of our system’s design. QOS comprises a layered architecture that consists of two main components: the QOS compiler (top) and the QOS runtime (bottom), which we detail next.
Qernel Abstraction. QOS implements a wide range of mechanisms with different abstraction requirements, from the compilation to the execution runtime level. To enable the composability of these mechanisms in a unified architecture, we propose the Qernel abstraction that acts as a common denominator for the QOS mechanisms to apply their policies.
QOS Compiler. We propose the QOS compiler (Figure 5, top), a modular, extensible, and composable compiler infrastructure. It comprises three stages: (1) The frontend of the compiler, the analyzer (§ 5.3), accepts quantum circuits and lifts them to the Qernel abstraction, generates the intermediate representation (IR), and performs IR analysis passes to generate the IR static properties required by the next stages. (2) The middle-end, the optimizer (§ 6), is an extensible and composable set of optimization passes that leverages the IR and static properties to improve the execution fidelity of the quantum circuits with manageable overheads. (3) The backend, the virtualizer (§ 7), compiles the optimized Qernels for the target QPUs, similar to classical target code generation.
QOS Runtime. We propose the QOS runtime (Figure 5, bottom), a system that abstracts away the underlying heterogeneity and balances the tradeoff between the conflicting objectives of the cloud operator (resource efficiency) and the users (high fidelity and low waiting times). The runtime comprises four components: (1) The estimator predicts the fidelity of executing the optimized Qernels to guide scheduling decisions. (2) The multi-programmer, given the estimations, bundles low utilization Qernels to increase QPU utilization. (3) The scheduler multiplexes and runs the Qernels across space and time with the objective to maximize fidelity and minimize waiting times. Finally, (4) the knitter post-processes the Qernel execution results to return the final result to the user.
Figure 6. Compiler Frontend: Analyzer (§ 5). (a) An example quantum circuit of 4 qubits and 5 gates. (b) The Qernel intermediate representation (QIR) (§ 5.1). (c) The refined QIR (§5.1). (d) The Qernel’s static and dynamic properties (§5.2). During compilation, the dynamic properties are void, but they are initialized and used during the runtime.
4.2. Execution Workflow
First, users submit a circuit along with their optimization target and budget 1. The former represents the desired post-compilation circuit size, and the latter quantifies the additional overheads the user is willing to pay. The compiler’s frontend lifts the circuit to the Qernel abstraction and generates the IR and its static properties 2. The middle-end optimizes the Qernels through a modular set of passes 3, then the backend generates the target QPU-optimized Qernels and submits them to the QOS runtime 4. The estimator predicts the fidelity of running the Qernel(s) on the QPUs to guide scheduling 5. The multi-programmer bundles Qernels with low utilization and sends them to the scheduler 6. The scheduler assigns and runs the bundled Qernels, optimizing for maximal fidelity and minimal waiting times 7. After the execution, the bundled results are retrieved by the multi-programmer 8 to be unbundled into separate results and are sent to the knitter 9. Finally, the knitter post-processes the separated results pass and returns them to the user 10.
5. Compiler Frontend: Qernel & Analyzer
5.1. The Qernel Abstraction
The Qernel is the unified abstraction acting as a common denominator for the QOS components. Specifically, a Qernel contains (1) the graph-based IR used by the QOS compiler and (2) the Qernel properties, which comprises static IR properties and dynamic properties used by the QOS runtime.
Qernel Intermediate Representation (QIR). Existing optimization techniques operate at the gate level, analogous to the instruction level in the classical domain. Therefore, we propose a graph-based Qernel Intermediate Representation (QIR) that captures the control flow of a quantum program, similar to the control flow graph of classical programs. By traversing the QIR, the compiler can identify important optimization opportunities, such as pairs of gates that cancel each other (like dead code elimination), gate dependencies (useful for gate scheduling, similar to instruction scheduling), or opportunities to remove hotspot gates (gates that contribute to noise errors in the computation). An example QIR is shown in Figure 6 (b), where the quantum circuit consisting of four qubits and five gates is lifted to the QIR.
Formally, a QIR is a directed acyclic graph (DAG) $G=(V,E)$, where $V$ is the set of gates and every edge $e_{i}\in E$ is the qubit the gate acts on. The edges’ directions reflect dependencies $D=\{(V_{i},V_{j})\in V\times V\}$ between gates, i.e., $V_{i}$ must be scheduled before $V_{j}$. To identify hotspot nodes, we compute the degree of a node $deg(V_{i})$, which reflects the number of control flow paths the gate $V_{i}$ is part of. In the example of Figure 6, the QIR reveals four layers of gates: $l_{1}:(g_{0},g_{1})$, $l_{2}:g_{2}$, $l_{3}:(g_{3},g_{4})$, and $l_{4}:M$, which means that they have to be scheduled in this order, and the pairs in the same layers are susceptible to crosstalk noise (§ 2). Finally, the gate $g_{2}$ is a hotspot node since its degree is 4, the highest in $V$ (the measurements $M$ are terminal nodes and do not count).
Refined QIR. Various optimizations require a simplified representation that captures only the connectivity between qubits. For instance, the connectivity structure might reveal hotspot qubits (Ayanzadeh et al., 2023) that can be removed for fidelity improvement or opportunities for circuit cutting (§ 6). Figure 6 (c) shows the refined QIR, where we can see that $q_{1}$ and $q_{2}$ are only connected by a single gate, and removing it would split the circuit into two smaller circuits. Formally, a refined QIR is an acyclic, undirected, and weighted graph $G=(V,E)$, where $V$ is the set of qubits and every edge $e_{i}\in E$ between two qubits $V_{j},V_{k}$ has a weight $w_{i}\in\mathbb{N}$ that represents the number of gates that act on $V_{j}$ and $V_{k}$.
5.2. Qernel Properties
For applying the diverse set of its mechanisms, QOS requires data structures that keep up-to-date information about the quantum programs. Such information includes (1) the static properties, which are useful for the compiler, and (2) dynamic properties, which are useful for the runtime.
Static Properties. Apart from the IR, optimization passes leverage circuit properties to be more efficient and effective. The properties include the circuit’s size (number of qubits), depth, non-local gates and their types, the number of measurements, and others (Figure 6 (d)). Additionally, we include the features vectors defined in (Tomesh et al., 2022) since they are potentially useful for heuristic-based optimizations or regression-based prediction models (Quetschlich et al., 2023).
Dynamic Properties. The Qernel also contains dynamic properties required by the QOS runtime. These include the Qernel’s execution status (done, failed, running, scheduled), the estimator’s output, i.e., fidelity estimations (§ 8.1), and the final post-processed results (Figure 6 (d)).
Figure 7. Optimization workflow (§ 6). The initial refined QIR has 7 qubits and 12 gates. (a) The HotSpotNodePass identifies $q_{3}$ as a hotspot node with a degree of 6. (b) We optimize the refined QIR with $s=2,b=3$ by applying the QubitFreezingPass which reduces the qubit and gate counts to 6. (c) We apply the GateCuttingPass to further remove 2 gates. This gives two fragments of 3 qubits each. (d) We have depleted the budget, so we use the QubitReusePass to achieve the goal of $s=2$.
5.3. Frontend: Analyzer
To discover and leverage optimization opportunities, we first need to perform circuit analysis, similar to classical program analysis. The analyzer transforms a quantum circuit into a Qernel and comprises an extensible set of passes that generate the QIR and static properties of the Qernel, which are then used by the optimizer and subsequently by the runtime.
QIR Transformation Pass. The first step in program analysis and optimization is to generate the QIR, implemented by the QIR transformation pass. Figure 6 (a)-(b) shows the generation of the QIR for an example quantum circuit. To generate the QIR, the pass iterates over each logical qubit of the circuit and each gate acting on that qubit. For each such gate, it creates a QIR vertex $g_{i}$ and sets the qubits $q_{j},q_{k}$ the gate acts on as the edges $e_{j},e_{k}$ of $g_{i}$. By convention, the direction of the edge follows the direction from the control to the target qubits. When reaching measurement operations, it simply adds the terminal nodes $M$. This process is repeated until all circuit qubits and gates are covered.
QIR Refinement Pass. To generate the refined QIR (Figure 6 (c)), we implement a transformation pass that traverses the QIR in a depth-first manner (breadth-first is equivalent). For each QIR node visited, i.e., a vertex $g_{i}$ with a pair of edges $q_{j},q_{k}$, it checks the current refined QIR for existing nodes with the same name. If true, it increments the weight of the edges between the nodes by one. Otherwise, it adds $q_{j}$ or $q_{k}$ or both as new nodes in the refined QIR and connects them with a weight of one.
Analysis Passes. We implement passes that analyze the QIR to identify optimization opportunities, such as gate dependencies (DependencyGraphPass), hotspot nodes (HotspotNodePass), and graph isomorphism (IsIsomorphicPass). We also implement properties passes that traverse the QIR to generate the Qernel static properties (§ 5.2) and comprise the BasicAnalysisPass and SupermarqFeaturesPass. Specifically, the former generates the key circuit properties while the latter computes the six feature vectors defined in (Tomesh et al., 2022), as explained in § 5.2. We show how the optimizer uses the information obtained from these passes in § 6.
6. Compiler Middle-end: Optimizer
QPUs comprise up to only a few 100s qubits, which sets the physical limit for circuit size and are noisy, which sets a practical limit to high fidelity execution (§ 3.1). To increase the scalability of circuits that run with high fidelity, we need a modular, extensible, and composable optimizer. Modular to support adding/removing optimization passes or changing their relative order, extensible to add new passes, and composable to chain the optimization improvements of the individual passes.
Such optimization passes are the circuit compaction techniques that, as the name suggests, reduce the circuit size, i.e., the number of qubits, rendering it executable on small QPUs and at the same time, also simplify the circuit structure, i.e., remove noisy gates. Notably, we can compose more than one of these techniques to achieve even better results. The compiler’s middle-end, the optimizer, is a composable pipeline of transformation passes that compact the QIR to increase the scalability of quantum circuits running with high execution fidelity.
Challenges. However, it is not trivial to implement such a pipeline. Currently, there is a plethora of individual compaction techniques that require their own sub-systems to operate, with no common infrastructure to compose them. Additionally, as we will show later, some techniques spawn an exponential number of sub-circuits (Table 1) and, after execution, require post-processing using classical hardware. We pose two questions: (1) How are the (exponentially) spawned circuits from different techniques handled? (2) How can we manage the tradeoff between fidelity improvement and exponential overheads from different techniques?
Our Approach. To this end, we design our optimizer with the goals of providing (1) a unified infrastructure for pluggable compaction mechanisms and (2) tunable knobs for configuring the tradeoff between overheads and performance improvement. For (1), we build and compose vastly different compaction techniques on the Qernel abstraction, specifically on the QIR and its refined form. For (2), we provide users with two knobs: the optimization budget (equivalent to optimization level) $b\in\mathbb{N}$ and the size to reach $s\in\mathbb{N}$, which denotes the desired post-optimization QIR size (number of qubits). Since all overheads are exponential and the exponent’s base does not make a practical difference, the single budget knob, $b$, suffices.
QIR Compaction Techniques. There are two main QIR compaction techniques: circuit divide-and-conquer and qubit reuse. In the former category, the (large) QIR is cut into smaller fragments that are executed on small QPUs, and the execution results are merged back to a single value. Circuit cutting and knitting (Peng et al., 2020; Mitarai and Fujii, 2021) belongs to this category. In our optimizer, we implement a pass that automatically cuts the QIR based on qubits (WireCuttingPass) and a pass that cuts the QIR based on gates (GateCuttingPass). To restrict the exponential overheads that scale with the number of cuts, we use the budget $b$ to cut up to $b$ times. At each cut location, the pass places a virtual gate, which must be later replaced with other gates that simulate the effects of the original (pre-cut) gate (§ 7.1).
We implement another circuit divide-and-conquer technique, also with exponential overheads, namely the QubitFreezingPass, which is limited to QAOA applications only (Ayanzadeh et al., 2023). At a high level, it removes qubits with significantly more noisy gates than other qubits, i.e., hotspot qubits, along with the gates. This reduces the number of physical qubits required by an underlying (small) QPU and greatly reduces the number of noisy gates. We use the same budget $b=3$ to remove the nodes with the highest degrees to restrict the exponential overheads.
Lastly, in the qubit reuse category, we implement the QubitReusePass, which “compacts” multiple logical qubits into one to reduce the QIR’s size (Hua et al., 2023; DeCross et al., 2022; Sadeghi et al., 2022). This process, however, increases the QIR’s depth; therefore, the tradeoff, in this case, is between QIR size and depth. To restrict the depth increase, we use qubit reuse as a last resort to achieve the user’s size requirement ($s$) or to render the Qernel executable by at least one QPU in the system.
Optimization Workflow. Figure 7 shows the default optimization workflow for the refined QIR of a QAOA circuit (§ 2.1) with 7 qubits and 12 gates. The optimizer aims to achieve a maximum QIR size $s=2$ with an allowed budget $b=3$. To achieve this, it takes the following steps:
Step 1: The optimizer calls the HotspotNodePass pass on the refined QIR to find a hotspot node. The pass identifies $q_{3}$ as a hotspot with a degree of 6 gates (Figure 7, (a)).
Step 2: The optimizer applies the QubitFreezingPass to remove $q_{3}$ and its gates. The new refined QIR size is 6 qubits with 6 gates. Then, it updates the budget to $b=b-m$, where $m=1$ is the number of qubits frozen (Figure 7, (b)).
Step 3: The optimizer applies either gate or wire cutting. To do so, it first computes the expected cost $c_{gate},c_{wire}$, respectively, to achieve a circuit of size $s=2$, and selects the one with the lowest cost $c_{min}=min(c_{gate},c_{wire})$. In this case, the cost for gate cutting is lower, so it applies the GateCuttingPass on 2 gates and updates the budget to $b=0$. The new refned QIR size is two fragments of 3 qubits and 4 gates each (Figure 7, (c)).
Step 4: Since $b=0$ but $s_{QIR}>s$, the optimizer applies the QubitReusePass to achieve $s=2$. The pass identifies qubits $q_{0},q_{1}$ as reusable and applies measurement and reset to them. The final refined QIR now has two fragments of $s=2$ and 4 gates (Figure 7, (d)).
The final optimizer’s output is a Qernel with a 42.8% smaller size and 66% less noisy gates. Note that each of the above passes alone wouldn’t achieve this result.
Figure 8. (a) Compiler Backend: Virtualizer (§ 7) and (b) QOS Runtime: Knitter (§ 8.4). (a) The Virtualizer consists of two stages: (1) Instantiation, which transforms the optimized Qernel to instantiated sub-Qernels (ISQs). This process generates $O(6^{k})$ ISQs for $k$ cuts (here $k=2$). (2) Target QPUs transpilation, where the ISQs are transpiled to every QPU architecture or every QPU in general. (b) The knitter consists of two stages as well: (1) The map phase and (2) the reduce phase.
7. Compiler Backend: Virtualizer
The backend stage of the QOS compiler, the virtualizer, generates the final executable Qernels for the underlying runtime, similar to classical compilers that generate the target code. The virtualizer consists of two stages: (1) the instantiation, which replaces the virtual gates from the cutting optimization passes with the gates that simulate the original ones, and (2) target QPUs transpilation, which translates the high-level gates to the physical QPU gates and performs mapping and routing, as explained in § 2.2
Optimization Pass # ISQs Generated Post-processing
Wire Cutting $O(4^{k}-8^{k})$ $O(4^{k}-8^{k})$
Gate Cutting $O(6^{k})$ $O(6^{k})$
Qubit Freezing $O(2^{k})$ $O(1)$
Table 1. QOS Virtualizer (§ 7.1). Number of instantiated sub-Qernels (ISQs) generated and post-processing complexity for each optimization pass, as a function of the number of cuts $k$.
7.1. Instantiation
The circuit cutting and knitting passes we describe in § 6 cut a large Qernel into sub-Qernels by analyzing the QIR, identifying optimal cut locations, and then placing virtual gates there. However, to run the sub-Qernels, we must replace the virtual gates with a combination of 1-qubit gates that achieves the same computation as the original Qernel. The mapping between virtual and 1-qubit gates depends on the chosen cutting strategy, i.e., gate or wire cutting (§ 6). We refer to this process as instantiation.
The instantiation stage takes as input the optimized sub-Qernels with virtual gates and outputs instantiated sub-Qernels (ISQs) with the 1-qubit gates required to execute them. Similar to the general cutting approach, we implement a generic instantiation mechanism for supporting pluggable mappings from virtual to 1-qubit gates. The mappings depend on the cutting technique (i.e., gate and wire cutting require different mappings) but can also differ for the same technique. For instance, virtual gates from gate cutting can be mapped to different sets of 1-qubit gates that might be more optimal for specific QPU technologies.
By replacing a single virtual gate with multiple 1-qubit gates, the mapping function generates multiple ISQs that differ only by the 1-qubit gate. Then, replacing the next virtual gate in each ISQ generates even more copies, which leads to an exponential number of ISQs. Figure 8 (a) shows a Qernel optimized using two gate cuts. The red boxes are the two virtual gates that must be replaced with 1-qubit gates. In this example, the mapping function will replace the first virtual gate with six 1-qubit gates, creating six ISQs. For the next and final virtual gate, each of the six ISQs will produce six more ISQs, totaling 36 ISQs. Generally, in QOS, the exact overheads are $O(2^{k}-8^{k})$ for $k$ cuts for our optimization passes (Table 1).
7.2. Target QPUs Transpilation
Following instantiation, the ISQs must be transpiled (§ 2.2) to the target QPUs to be sent to the runtime for scheduling and execution. Since the number of ISQs might be large, depending on the optimization budget $b$ used at the compiler middle-end (§ 6), we offer two transpilation modes that differ in granularity and overheads: (1) the coarse-grain per QPU architecture and (2) the fine-grain per QPU. We show evaluation results for their transpilation overheads in § 9.2.
Per QPU-Architecture. In the first mode, we transpile each ISQ for every type of QPU architecture available in the system. This coarse-grain approach bounds the transpilation overheads because typical quantum cloud providers have a limited number of architectures, e.g., up to five (ibmQuantum, [n. d.]). Since this mode does not scale with the number of QPUs, our experimentation shows that it is suitable for values of budget $b\geq 5$, which generate $10^{4}-10^{5}$ ISQs.
Per-QPU. In the second mode, we transpile each ISQ to each available QPU in the system. This will enable the runtime components to make fine-grained decisions about the fidelity of running the ISQ on any QPU since they will have the exact noise information of this ISQ-QPU pair. The overheads are still bound since QPUs are constant in quantity in commercial clouds, e.g., up to 30 (ibmQuantum, [n. d.]), in contrast to classical clouds that scale to thousands of classical nodes. Our experimentation showed that this transpilation mode is viable for $b<5$.
8. QOS Runtime
The QOS runtime (Figure 5, bottom) schedules and executes Qernels across space and time in a scalable manner to achieve the user’s goals, i.e., higher fidelity and lower waiting times, and the cloud operator’s goals, i.e., resource efficiency. It comprises four components, which we detail next. For simplicity, we use the general term Qernel throughout this Section.
8.1. Estimator
The estimator is responsible for predicting the fidelity of a given Qernel on the underlying QPUs without executing the Qernel. This prediction will be the leading decision factor for the scheduler when assigning the Qernel to a QPU. To achieve this, it computes a score for each Qernel-QPU assignment that captures the potential fidelity of that assignment and then uses the scores to rank the assignments. The estimator supports configurable scoring policies that consider (1) the Qernels’ properties generated from the compiler and (2) the QPUs’ calibration data, which are available to quantum cloud providers since they perform the calibration cycles.
For (1), important properties include the number and types of gates, depth, and the number of measurements (§ 5.2). For (2), recall that QPUs are characterized by calibration data that describe the exact error rates of the QPU for that calibration cycle (§ 2.2), specifically, the individual qubit readout errors, the individual gate errors, and the $T2$ coherence times. In this work, we implement two scoring policies: a numerical approach for fine-grained control over the estimations and a regression model approach for abstracting away the complexity of estimation.
Numerical Cost Policy. This policy estimates execution fidelity by leveraging the target-QPU transpilation output of the compiler backend (§ 7). Target transpilation enables fine-grained fidelity estimation by producing the mapping between logical and physical qubits and the gate (instruction) schedule. The mapping captures the expected readout and gate errors, while the gate schedule captures the order and exact timing that the gates will be applied on the qubits, which reveals the hardware decoherence and crosstalk errors, as explained in § 2.2.
Formally, for each qubit $q_{i}$ the readout error is $e_{r(i)}$, for each gate $g_{j}$ the error is $e_{g(j)}$, and the decoherence error is $e_{d(t)}=1-e^{-t/T2_{i}}$, where $t$ is the idle time of the qubit $q_{i}$ (no gates act on it (Das et al., 2021b)) and $T2$ is the decoherence time of $q_{i}$. The crosstalk error between gates $g_{k}$ and $g_{l}$ is $e_{ct(k,l)}$. Putting it all together, the final fidelity score is computed as follows: $fid=1-\prod_{i=0}^{N}e_{r(i)}e_{d(i)}\prod_{j=0}^{M}e_{g(j)}\prod_{j=0,k=0}^{M\times M}e_{ct(j,k)}$, where $N$ is the circuit’s number of qubits and $M$ is the number of gates. Since all hardware error information is known at-priori, and quantum errors accumulate multiplicatively, this policy produces high-accuracy estimations, as we show in § 9.3.
Figure 9. QOS multi-programmer example workflow (§ 8.2). (a) Use the estimator’s output to find Qernels with the same best QPU, (b) compute their independent utilization, and (c) compute their compatibility score. If compatible, (d) check for layout overlap, and (e) apply the appropriate multi-programming policy.
Regression Model Policy. As discussed in § 2.2, QPU noise errors are accurately measured at each calibration cycle, and their impact on fidelity during quantum computation can be described mathematically. Therefore, we can train a regression model to predict the fidelity of a transpiled Qernel on a possible QPU using the QPU’s calibration data and the Qernel’s static properties as features (Quetschlich et al., 2023; Ravi et al., 2021b). Specifically, we use the aforementioned errors we defined in the numerical cost policy as QPU features and the static properties (§ 5.2) as Qernel features. Even simple regression models such as linear regression achieve high prediction accuracy, up to 99%. This policy is simple to use without detailed knowledge of the relationship between errors. However, in QOS, we use the numerical cost policy by default for estimation to have a clear understanding and full control of the process.
8.2. Multi-programmer
The size of quantum programs that run with high fidelity is small, leading to QPU underutilization (§ 3.3). To increase QPU utilization, QOS multi-programs two or more Qernels, potentially from different users, to run on the same QPU. We refer to this multi-programming as bundling the Qernels together. However, trivially bundling Qernels together will deteriorate fidelity because qubits interfere with each other via crosstalk errors (§ 2.2). On top of that, bundled Qernels that run for unequal durations do not necessarily increase utilization since QPU effective utilization is measured in space (number of QPU qubits allocated) and time (time qubits are performing actual computation).
For example, a 10-qubit Qernel $Q_{0}$ running on a 20-qubit QPU gives 50% spatial utilization. However, assume that $Q_{0}$ runs 3$\times$ longer than a 10-qubit Qernel $Q_{1}$. During $\frac{2}{3}$ of $Q_{0}$’s runtime, the qubits allocated to $Q_{1}$ will be idle, decreasing the effective utilization to only 66%. Recall that it is impossible to schedule more Qernels during $Q_{0}$’s runtime, unlike in a typical CPU (§ 2.2).
To minimize the fidelity impact and maximize the effective utilization of multi-programming, we utilize configurable Qernel compatibility functions that quantify how well-suited are two Qernels to run together.
Qernel Compatibility Functions. Compatibility functions measure the crosstalk errors and the effective utilization of a Qernel pair by considering the Qernels’ static properties (§ 5.2). To measure crosstalk effects, we identify pairs of 2-qubit gates that run in parallel during Qernel execution. To quantify this without running the bundled Qernels, we use the entanglement ratio and parallelism Qernel static properties, where higher values indicate a higher chance for crosstalk errors (Tomesh et al., 2022). Intuitively, the entanglement ratio captures the proportion of 2-qubit gates over all gates, and parallelism captures how many gates run in parallel per time unit, on average.
To measure the spatial dimension of effective utilization, it suffices to compute the ratio of allocated QPU qubits over the number of QPU qubits. To measure the temporal dimension, we compare the relative duration between two Qernels. The depth static property reflects the longest chain of gates that will be executed; therefore, it measures the Qernel’s duration. More technically, we define effective utilization as $u_{eff}=\frac{N_{C_{max}}}{N_{QPU}}*100+\sum_{n=1}^{k}\frac{D_{n}}{D_{max}}*\frac{N_{C_{n}}}{N_{QPU}}*100$, where $N_{C_{max}},N_{QPU}$ are the number of qubits of the longest Qernel and the QPU, respectively, $k$ is the number bundled Qernels excluding the longest Qernel, and $D$ is the depth of the Qernel.
To put everything together formally, we score a possible Qernel pair as follows: $qc=\alpha~{}u_{eff}+\beta~{}ER_{b}+\gamma~{}PA_{b}\mapsto[0,1]$, where higher is better, $\alpha+\beta+\gamma=1$, and $c$ denotes bundled, i.e., $E_{b}$ is the entanglement ratio of the bundled Qernels. The four variables are tunable to give priorities on different objectives, e.g., prioritize effective utilization or minimize crosstalk. After experimenting and fine-tuning, we found that $\alpha=0.25,\beta=0.25,\gamma=0.5$, and $qc\geq 0.75$ gives balanced results, as we show in § 9.4.
Figure 9 shows an example workflow. The multi-programmer receives three Qernels with three estimations each and identifies $Qernel~{}0$ and $Qernel~{}2$ as a possible pair since their best QPU is the same ($QPU5$) (a). It computes their independent utilization, which is 31% and 37%, respectively, and the combined utilization is under 100% (b). It computes the compatibility score that surpasses the threshold ($0.9>0.75$) (c). Next, we detail our multi-programming policies.
Multi-programming Policies. QOS supports pluggable multi-programming policies for maximizing effective utilization or minimizing fidelity penalties. In this work, we implement two multi-programming policies; the first is the fast path multi-programming, where we can immediately bundle two Qernels if there is no conflict between them, while the second requires re-compilation and re-estimation.
Restrict Policy. The restrict policy uses the target QPU transpilation output to bundle Qernels if there is no overlap in their layouts. Practically, this means that for Qernels $Q_{0}$ and $Q_{1}$, their logical qubits are mapped to disjoint sets of physical qubits on the QPUs. In that case, the policy bundles the Qernels together, and fidelity loss is minimized through the aforementioned compatibility score.
Re-evaluation Policy. This policy is the fallback of the restrict policy. If the Qernel layouts overlap, the two Qernels are transpiled again for the target QPU, and their new fidelity is estimated. If the new fidelity is lower up to a fixed $\epsilon>0$ value compared to the original fidelities, the bundling is maintained. Otherwise, the multi-programmer selects the next most compatible Qernel pair.
Figure 9 (d) shows the check for layout overlap. In this example, yellow qubits belong to $Qernel~{}0$ and green to $Qernel~{}1$. On the left, there is no overlap, while on the right, the red qubit is shared between the Qernels. (e) We apply the respective policy (in this example, re-evaluation).
8.3. Scheduler
Figure 10. QOS Compiler (§ 9.2). Impact of the QOS compiler on the circuit depth and the number of CNOTs. The circuits are optimized using budget $b=3$, and we compare against Qiskit (red horizontal line), FrozenQubits (Ayanzadeh et al., 2023) and CutQC (Tang et al., 2021b). There is an average $46\%$, $38.6\%$, and $29.4\%$ reduction in circuit depth, and $70.5\%$, $66\%$ and $56.6\%$ reduction in the number of CNOTs, respectively.
Scheduling quantum programs involves fundamental tradeoffs between conflicting objectives; specifically, users want maximal fidelity and minimal waiting times. However, to maximize fidelity, most programs must run on the same subset of QPUs that perform best in a given calibration cycle (§ 3.2). This will lead to large and growing queues on these QPUs, hence long waiting times for the users.
Our scheduler assigns and runs Qernels across space (which QPUs) and time (when) and supports pluggable policies for managing the aforementioned tradeoffs, prioritizing maximal fidelity, minimal waiting times, or a balanced approach. The scheduler assigns Qernels to QPUs based on the fidelity estimations provided by the estimator and the execution time estimations, which we detail next.
Execution Time Estimation. To optimize for minimal waiting times, the scheduler must first estimate each Qernel’s execution time and then aggregate the execution time estimations in each QPU’s queue to compute the total waiting times. To estimate the execution time, we iterate the longest path of the QIR of a Qernel (§ 5.1) that corresponds to the longest-duration gate chain and thus defines the Qernel’s execution time. By summing the gate durations of each node in the longest path, we get the Qernel’s total execution time.
Formula-Based Policy. Optimizing for conflicting objectives involves comparing two possible solutions (e.g., maximal fidelity vs. minimal waiting times). In the formula-based policy, we use a simple scoring formula (Equation 1) to compare and select between two possible assignments. This formula factors fidelity, waiting time, and utilization to determine which assignment is better, given priorities. The parameters are as follows: $f_{i}$: fidelity of the estimation result $i$, $t_{i}$: waiting time for the QPU from estimation result $i$, $u_{i}$: utilization of the QPU for estimation result $i$, $c\in(0,1)$: a system-defined constant that weighs the fidelity difference between estimations and finally, $\beta$: a system-defined constant acting as a weighting factor for utilization difference, balancing system throughput and fidelity. By selecting higher $c$, the system prioritizes fidelity over waiting times, and vice versa, and by selecting higher $\beta$ the system prioritizes utilization over fidelity, and vice versa. By default, $c=\beta=0.5$, which aims for balanced fidelity, waiting times, and utilization.
(1) $Score=c\frac{f_{2}-f_{1}}{f_{1}}-(1-c)\frac{t_{2}-t_{1}}{t_{1}}+\beta\frac{u_{2}-u_{2}}{u_{1}}$
Genetic Algorithm Policy. Genetic algorithms excel at optimizing for conflicting objectives by efficiently searching over vast search spaces, and for that, they can be used in the context of QOS. We formulate a multi-objective optimization problem with the conflicting objectives of fidelity vs. waiting times and use the NSGA-II genetic algorithm (Deb et al., 2002) to solve it. The algorithm creates a Pareto front of possible solutions (schedules), each achieving a different combination of average fidelity and average waiting times. Then, to select one of those schedules, we use the formula described by Equation 1 to score each schedule and select the schedule with the highest score.
8.4. Knitter
Following scheduling and execution, the QOS runtime collects the results that are part of the initial circuit submitted by the user. Recall that the circuit is lifted to the QIR (§ 5.3), then optimized through divide-and-conquer techniques that place virtual gates inside the QIR (§ 6), and finally instantiated to replace the virtual gates with 1-qubit gates (§ 7.1). The instantiation process generates up to $O(8^{k})$ instantiated sub-Qernels (ISQs) for a single initial optimized Qernel (Figure 8 (a), Table 1). Finally, the ISQs are bundled with other ISQs, possibly from other users, for increased utilization (§ 8.2).
Therefore, to compute and return the final result to the user is not trivial; we must first unbundle the results from multi-programming and then merge the results from ISQs to the original Qernel, a process called knitting. The structure of the ISQs and their respective results resembles a tree structure, where the leaf nodes are up to $O(8^{k})$ results, and the root node is the final result. Therefore, we adopt the map-reduce pattern to perform knitting.
Unbundling for Multi-programming. The results first pass through the multi-programmer to be unbundled. To do this, the multi-programmer keeps a record that maps the initial (solo) Qernel IDs to the new, bundled Qernel ID, as well as the Qernels’ sizes. Therefore, when receiving a new result from a Qernel with an ID $i$, it scans the record to find an entry $i$, and if found, it splits the probability distribution bitstrings (§ 2) into two parts: the left-most and the right-most bits based on the Qernel sizes. Then, it forwards the unbundled results to the knitter for the map-reduce phases.
The Map Phase. To efficiently process a large number of results (up to $O(8^{k})$), we follow a divide-and-conquer approach. Specifically, we split the results into $k$ equal sizes and distribute them to $k$ classical nodes to be processed in parallel (Figure 8 (b), step (1)). We parallelize across $k$ to increase data locality and reduce communication overheads since all results for each of the $k$ cuts will be in the memory of the same node. Locally, each node performs tensor product ($\bigotimes$) operations on the probability distributions, which are parallelizable across the node’s threads. If available in the node, QOS leverages GPUs or TPUs to accelerate the tensor products. Following this process, the $k$ nodes output $k$ intermediate results, ready to be reduced into a single result.
The Reduce Phase. QOS selects any of the $k$ nodes to perform the reduce step. The rest of the nodes send the intermediate results to this node, which performs a thread-parallel sum of $k$ results. Equivalent to the map phase, the parallel sum can also be executed on GPUs. This produces the final output to be returned to the user (Figure 8 (b), step (2)).
9. Evaluation
9.1. Experimental Methodology
Experimental Setup. We conduct two types of experiments: (1) classical tasks, such as circuit transpilation and trace-based simulations, and quantum tasks (2), which run on real QPUs for measuring the circuits’ fidelities.
Figure 11. QOS Compiler (§ 9.2). Impact of the QOS compiler on the circuit fidelity against Qiskit (qiskit-transpiler, [n. d.]), CutQC (Tang et al., 2021b), and FrozenQubits (Ayanzadeh et al., 2023). The circuits are optimized using budget $b=3$. There is a mean 2.6×2.6$\times$2.6 ×, 1.6×1.6$\times$1.6 ×, and $1\times1$ improvement for 12-qubit circuits, respectively. There is a 456.5×456.5$\times$456.5 ×, 7.6×7.6$\times$7.6 ×, and 1.67×1.67$\times$1.67 × improvement for circuits of 24 qubits, respectively.
For (1), we use a server with a 64-core AMD EPYC 7713P processor and 512 GB ECC memory. For (2), we conduct our experiments on IBM Falcon r5.11 QPUs. Unless otherwise noted, we use the IBM Kolkata 27-qubit QPU.
Framework and Configuration. We use the Qiskit (Qiskit contributors, 2023) Python SDK for compiling quantum circuits and running simulations. We compile quantum circuits with the highest optimization level (3) and run with 8192 shots. Each data point presented in the figures is the median of five runs.
Benchmarks. We study QOS on a set of circuits used in state-of-the-art NISQ algorithms, adopted from the 3 benchmark suits of Supermarq (Tomesh et al., 2022), MQT-Bench (Quetschlich et al., 2022) and QASM-Bench (Li et al., 2023). The algorithms’ circuits can be scaled by the number of qubits and depth. Specifically. We study 9 benchmarks: GHZ, W-State, Bernstein Vazirani (BV), Hamiltonian Simulation (HS-$t$), Quantum-enhanced Support Vector Machine (QSVM), Two Local Ansatz (TL-$n$), Variational Quantum Eigensolver (VQE-$n$), and Approximate Optimization Algorithm (QAOA-R/P), these benchmarks cover a wide range of relevant criteria for evaluating QOS.
For the TL and VQE circuits, we use circular and linear entanglement, respectively. The HS, VQE, and TL benchmarks are scalable by their circuit depth with the number of time-steps $t$ and layers in the $ansatz$ $n$. The QAOA-R/P circuits are initialized using regular/power-law graphs, respectively, with degree $d\in\{1,3\}$.
Metrics. We evaluate the following metrics:
  • Fidelity: We use the Hellinger fidelity as a measure of how close a noisy result is to the desired ground truth of a quantum circuit (Hellinger, 1909; fidelity-qiskit, [n. d.]). The Hellinger fidelity is calculated as $Fidelity(P_{ideal},P_{noisy})=\left(1-H\left(P_{ideal},P_{noisy}\right)^{2}\right)^{2}\mapsto[0,1]$, where $H$ is the Hellinger distance between two probability distributions, and $P_{ideal},P_{noisy}$ are the ideal and noisy probability distributions, respectively.
  • Circuit Properties: Number of CNOT gates and depth. When a Qernel contains more than one sub-Qernel, we use the sub-Qernel with the maximum depth, amount of CNOTs, or an average of these two properties.
  • Waiting Time: The time a circuit spends in a QPU’s queue, waiting for execution, in seconds.
  • Classical Overhead: The optimization and post-processing overheads (§ 7) of the QOS compiler vs. Qiskit’s transpiler (qiskit-transpiler, [n. d.]).
  • Quantum Overhead: The number of additional quantum circuits we need to execute per original quantum circuit.
Baselines. We evaluate the QOS compiler against Qiskit v0.41, CutQC (Tang et al., 2021b) and FrozenQubits (Ayanzadeh et al., 2023). QOS’s multi-programmer is evaluated against (Das et al., 2019). Regarding QOS scheduler, to the best of your knowledge, (Ravi et al., 2021b) is the only peer-reviewed quantum scheduler, but it doesn’t provide source code or enough technical details to faithfully implement it.
9.2. QOS Compiler
Figure 12. QOS Compiler (§ 9.2) and QOS Estimator (§ 8.1). (a) Compiler: classical and quantum overheads and fidelity improvement as a relative factor to Qiskit. For 24 qubits, the improvement outweighs the overheads. (b) Compiler: scalability to a large, hypothetical 1000-qubit QPU. Any budget $b>0$ achieves higher quality results than using no optimizations. (c) Estimator’s performance: fidelity of IBM Auckland vs. the QPU automatically selected by the estimator.
RQ1: How well does the QOS compiler improve the fidelity of circuits that run on NISQ QPUs? We evaluate the performance of the QOS compiler w.r.t the post-optimization properties and fidelity of the circuits while also analyzing the classical and quantum costs of our approach.
Effect on the Circuit Depth and Number of CNOTs. In Figure 10, we show the performance of the QOS compiler on the circuits’ depth and number of CNOTs, where we plot the relative difference in post-optimization circuit depth and the number of CNOTs between Qiskit (the red horizontal line) and FrozenQubits (Ayanzadeh et al., 2023), CutQC (Tang et al., 2021b), and the QOS compiler. Figures 10 (a) and (c) show that the circuit depth decreases by $46\%$, $38.6\%$, and $29.4\%$, respectively. Figures 10 (b) and (d) show that the number of CNOTs decreases by $70.5\%$, $66\%$, and $56.6\%$, respectively. The improvement in both metrics against the baselines is attributed to the composability of our compiler; the combined effect of circuit compactions (§ 6) achieves better results than standalone techniques.
Impact on Fidelity. Figure 11 shows the QOS-optimized circuits’ fidelity against Qiskit (qiskit-transpiler, [n. d.]), CutQC (Tang et al., 2021b), and FrozenQubits (Ayanzadeh et al., 2023). The results show a mean 2.6×2.6$\times$2.6 ×, 1.6×1.6$\times$1.6 ×, and $1\times1$ improvement for 12-qubit circuits, respectively, and a 456.5×456.5$\times$456.5 ×, 7.6×7.6$\times$7.6 ×, and 1.67×1.67$\times$1.67 × improvement for circuits of 24 qubits, respectively. The fidelity improvement is a consequence of lower circuit depths and fewer CNOTs, as shown in Figure 10.
Classical and Quantum Overheads. Figure 12 (a) shows the average classical and quantum overheads of the QOS compiler. The classical overhead is 16.6×16.6$\times$16.6 × and 2.5×2.5$\times$2.5 × for 12 and 24 qubits, respectively, and the quantum overhead is 31.3$\times$ and 12$\times$ for 12 and 24 qubits, respectively. However, fidelity improves by 2.6×2.6$\times$2.6 × and 456.5×456.5$\times$456.5 × for 12 and 24 qubits, respectively; therefore, for larger circuits, the fidelity improvement is worth the cost.
Scalability. To demonstrate that the QOS Compiler increases the scalability, we run the VQE-1 benchmark on a hypothetical 1000-qubit QPU with one-qubit gate errors of $10^{-4}$, two-qubit gate errors of $10^{-3}$, and measurement errors of $10^{-2}$. We optimize with budget $b\in\{0,1,4,8\}$ and report the estimated fidelity. Figure 12 (b) shows that all budget $b$ values improve the estimated fidelity, with a tradeoff of improvement vs. overheads.
RQ1 takeaway: The QOS compiler improves the properties of quantum circuits by $51\%$ on average, leading to an improvement in fidelity of 2.6–456.5$\times$, while incurring acceptable classical and quantum overheads.
9.3. Estimator
RQ2: How well does QOS’s estimator address spatial and temporal heterogeneities? We evaluate the estimator’s precision in selecting the top-performing QPU for each benchmark. We establish a baseline using the on-average best-performing machine every calibration day. On the day of the experiment, IBM Auckland was the best-performing machine (also with the highest number of pending jobs).
Estimator’s Accuracy. Figure 12 (c) shows the fidelity of the eight benchmarks when run on QPUs selected by the estimator versus when run on the IBM Auckland QPU. The QPU selected for the BV benchmark is Auckland; therefore, we omit this result. For the rest of the benchmarks, the IBM Sherbrooke and Brisbane QPUs were automatically selected. Interestingly, the fidelity is on par or even higher than IBM Auckland, except for only one benchmark, the QAOA-P1.
RQ2 takeaway: QOS’s estimator automatically identifies QPUs with higher fidelity than the current standard practice.
9.4. Multi-programmer
RQ3: How well does QOS’s multi-programmer increase QPU utilization with minimum fidelity penalties? We evaluate the impact of the multi-programmer on the fidelity of co-scheduled circuits for certain utilization thresholds.
Utilization vs. Fidelity. Figure 13 (a) shows the average fidelity of nine benchmarks with utilization of 30%, 60%, and 88%. The three bars represent: no multi-programming (No M/P) refers to large circuits that run solo, baseline multi-programming (Baseline M/P) refers to (Das et al., 2019), and QOS’s multiprogramming approach (QOS M/P). There is an average 9.6$\times$ improvement in fidelity compared to solo execution and an average 15% (1.15$\times$) improvement compared to the baseline.
Figure 13. QOS Multi-programmer (§ 9.4). (a) Impact of multi-programming on fidelity. There is a 9.6$\times$ and 1.15$\times$ improvement compared to no multi-programming and the baseline, respectively. (b) Effective utilization. There is 7.2% higher effective utilization on average. (c) Relative fidelity w.r.t. solo circuit execution. There is an average 9.6% drop in fidelity due to QOS’s multi-programming.
Effective Utilization. The results in Figure 13 (b) show that QOS achieves, on average, a 7.2% higher effective utilization (§ 8.2), with a maximum improvement of 10.1%.
Fidelity Penalty vs. Solo Execution. In Figure 13 (c), we evaluate the fidelity penalty of multiprogramming vs. solo circuit execution for utilization of 30%, 60%, and 88%. The fidelity loss is 2%, 9%, and 18%, respectively. The average fidelity loss is 9.6% compared to solo execution, which is in line with previous studies (Das et al., 2019; Liu and Dou, 2021). In the worst case (18%), the fidelity loss is caused by the restrictions in high-quality qubit allocations and the crosstalk errors.
Figure 14. QOS Scheduler (§ 9.5). (a) Formula-based scheduling policy: Average fidelity vs. average waiting times. A fidelity weight $c=0.7$ achieves 5$\sim$ 5$\times$ 5 × lower waiting time for only $\sim 2\%$ lower fidelity. (b) Genetic algorithm policy: it creates a Pareto front of schedules, where a fidelity weight $c=0.5$ achieves 2×2$\times$2 × lower waiting times for $\sim 4\%$ fidelity decrease. (c) QPU load as the total runtime of each QPU for the formula-based policy. The maximum load difference between any two QPUs is $15.2\%$.
RQ3 takeaway: The QOS multi-programmer improves fidelity by 1.15–9.6$\times$ and effective utilization by 7.2% compared to the baselines while incurring an acceptable fidelity penalty ($<10\%$) compared to solo execution.
9.5. Scheduler
RQ4: How well does QOS’s scheduler balance fidelity vs. waiting times and balance the load across QPUs? We evaluate our scheduler by generating a representative workload consisting of a dataset we collected during the development of QOS.
Dataset Collection. During our exploration of the motivational challenges (§ 3) and experimentation and evaluation of the QOS components and their policies, we collected a dataset of $70.000$ benchmark circuits and more than $7000$ job runs in the quantum cloud. We use this dataset to simulate representative workloads, as we detail next.
Workload Generation. To generate realistic workloads, we monitored all available QPUs on the IBM Quantum Cloud (ibmQuantum, [n. d.]) for ten days in November 2023 to estimate the hourly job arrival rate. The average hourly rate is 1500 jobs per hour and is the baseline system workload for our evaluation.
[…]
QPU Load Balancing. Figure 14 (c) shows the QPU load as the total runtime each QPU was active, in seconds, for the formula-based policy. All QPUs handle similar loads, with a maximum difference of $15.2\%$.
RQ4 takeaway: QOS scheduler balances the trade-off between waiting times and fidelity by reducing them 5$\times$ and only 2%, respectively, while balancing the load across QPUs.
10. Related work
Quantum optimization techniques can be categorized as (1) qubit mapping and routing (Li et al., 2019; Murali et al., 2019; Tannu and Qureshi, 2019c; Zhang et al., 2021; Molavi et al., 2022; Tannu and Qureshi, 2019a; Zulehner et al., 2019; Wille et al., 2019; Siraichi et al., 2018; Patel et al., 2022a; Patel and Tiwari, 2021; Baker et al., 2021), (2) instruction/pulse scheduling (Das et al., 2021b; Tripathi et al., 2022; Zlokapa and Gheorghiu, 2020; Smith et al., 2022; Murali et al., 2020; Shi et al., 2019a; Gokhale et al., 2020), (3) gate synthesis/decomposition (Shi et al., 2019a; Patel et al., 2022a; Litteken et al., 2023; Das et al., 2023; Xu et al., 2023; Patel et al., 2022b; Shi et al., 2019b), (4) execution post-processing and readout improvement (Bravyi et al., 2021; Maciejewski et al., 2020; Patel and Tiwari, 2020b; Das et al., 2021a; Patel and Tiwari, 2020a; Dangwal et al., 2023; Tannu and Qureshi, 2019b), and (5) circuit compaction (Mitarai and Fujii, 2021; Peng et al., 2020; Tang et al., 2021b; Bello et al., 2023; Ayanzadeh et al., 2023; Bichsel et al., 2020; Paradis et al., 2021; Ding et al., 2020; Paler et al., 2016; Hua et al., 2023). These techniques are implemented standalone without a compiler infrastructure and, thus, are not composable. Instead, the QOS compiler offers a powerful IR that enables incorporating such techniques in a composable manner.
Moreover, application-specific optimizations focus only on specific algorithms to enhance fidelity but lack generality (Alam et al., 2020; Lao and Browne, 2022; Gokhale et al., 2019b, a; Li et al., 2022; Stein et al., 2022; Resch et al., 2021; Hao et al., 2023; Tüysüz et al., 2023; Ravi et al., 2022a, 2023). In contrast, the QOS compiler is a generic approach applicable to all applications.
The state-of-the-art in quantum multi-programming is fairly limited (Das et al., 2019; Liu and Dou, 2021) and focuses solely on high-quality mapping, overlooking the systematic selection of compatible programs for utilization or fidelity. Notably, key optimizations from (Das et al., 2019) are integrated into the Qiskit transpiler workflow (qiskit-transpiler, [n. d.]), therefore, are already used by the QOS Virtualizer (§ 7.2).
Lastly, current quantum scheduling methods (Ravi et al., 2021b; Salm et al., 2022; Bhoumik et al., 2023) are limited because they (1) schedule circuits one at a time, (2) neglect QPU utilization, (3) lack fine control over waiting times versus fidelity, and (4) require manual input for final scheduling decisions. Work in the quantum cloud computing area (Karalekas et al., 2020; Leymann et al., 2020; Ravi et al., 2022b) and in quantum serverless (Nguyen et al., 2022; Garcia-Alonso et al., 2022; Grossi et al., 2021); describes quantum cloud characteristics or potential architectures, but QOS is the first end-to-end QPU management system.
11. Conclusion
We presented QOS, a system that composes cross-stack OS abstractions to address the challenges of quantum computing holistically. The synergy between compaction techniques, performance estimation, multi-programming, and scheduling systematically explores the tradeoff space associated with quantum. Specifically, QOS achieves up to 456.5×456.5$\times$456.5 × higher fidelity at a 12×12$\times$12 × overhead cost, up to 9.6×9.6$\times$9.6 × higher fidelity for a target utilization for $9.6\%$ lower fidelity than solo execution, and up to 5×5$\times$5 × lower waiting times for $2\%$ lower fidelity.
Contributions. Our main contributions include:
  1. (1)
    To our knowledge, QOS is the first attempt to combine circuit compaction with quantum resource management to tackle the challenges of QPUs holistically.
  2. (2)
    We leverage the QOS compiler infrastructure to compose optimizations that improve fidelity in a scalable manner, significantly outperforming their individual application (i.e., the current practice).
  3. (3)
    To our knowledge, we are the first to account for and improve both temporal and spatial QPU utilization when multi-programming quantum programs, while mitigating its associated fidelity penalties.
  4. (4)
    Our scheduler balances the inherent tradeoff between fidelity and waiting times, leading to better overall QoS.
Acknowledgements
We thank Karl Jansen and Stefan Kühn from the Center for Quantum Technology and Applications (CQTA)- Zeuthen for supporting this work by providing access to IBM quantum resources. We also thank Ahmed Darwish and Dmitry Lugovoy for their contributions to this work. Funded by the Bavarian State Ministry of Science and the Arts as part of the Munich Quantum Valley (MQV).